# Towards Variable-Length In-Network Caching

### v2

@@ 4 Implementation
Client-server application. We develop an open-loop application in C using NVIDIA Messaging Accelerator library (VMA) [3] that reduces host packet processing time. This enables the processing of packets in user space by implementing native RDMA verbs and bypassing kernel network stacks. The client application measures throughput and latency by generating requests. The time gap between consecutive requests follows an exponential distribution. The server application has multiple threads where each thread is pinned to a disjoint CPU core. To emulate multiple storage servers, we assign a partition per thread so that each thread acts as an independent server. This technique is used in NetCache [22] as well. In a similar vein, we implement a key-value store with TommyDS [1], a high-performance hash table library.
Switch. We implement the switch data plane in P416 [10] for Intel Tofino [6]. We use Intel P4 Studio SDE 9.7.0 to compile the switch data plane. Our prototype uses 10 match-action stages and 12.38% SRAM. The request table has 64K slots that can maintain up to 4K cached keys with a maximum queue size of 16 for each key. The controller is written in Python 3. In our prototype, we use an additional register array for the request table to store the timestamp of the request for latency measurement. The StarCache header also has 3 extra fields of 1-B Cached, 4-B Lat, and 1-B ParID. The first two fields are required to separately measure the latency of requests for cached and uncached items. The last field is to store the partition ID as we emulate multiple storage servers as dedicated threads in a physical node.
@@ 5.1 Methodology
Testbed setup. We build a cluster consisting of 8 nodes, which are connected by an APS BF6064X-T switch with the Intel Tofino 1 [6]. The servers are equipped with a 10-core CPU (Intel i5-12600K @ 3.7 Ghz), 32 GB of DDR5 memory, and a 100GbE NVIDIA CX-5 NIC. The servers run Ubuntu 22.04 LTS with Linux kernel 6.5.0. The 5 nodes act as clients to generate requests, and the remaining 3 nodes are used to emulate multiple storage servers.
@@
Workloads. We emulate 36 storage servers using 12 partitioned threads per node. We basically consider a workload with 10M key-value pairs whose key popularity skewness follows a Zipfian distribution with $\alpha=0.99$, since it is regarded as typical skewness [14, 8]. Although both key and value sizes impact whether an item is cacheable by NetCache, we use fixed 16-byte keys by default for simplicity. Instead, we represent the portion of cacheable items using different ratios between 64 bytes and 1024 bytes values. The 64-byte value represents a cacheable item value of NetCache. The 1024-byte value is a typical-sized item value in many workloads and is considered in many research papers [30, 31, 32, 18, 7, 27]. We use a bimodal distribution with 12% 64-byte and 88% 1024-byte values by considering the Cluster017 workload of Twitter [13]. In the workload, 12% of values are $\leq 128$ B and 87% of values are $>128$ and $\leq 1024$ B. 1% of values are between 1024 B and 1280 B, but we consider them 1024 B for simplicity. Most experiments are for read-only workloads like Cluster017 as we target read-intensive workloads.
Except for dynamic workloads, we try to preload 10K and 32 of static hottest items for NetCache and StarCache, respectively. 32 is a nearly optimal cache size for StarCache with the given workload. Since NetCache can cache only 12% of items in the workload, the actual number of cached items is 1.2K. We choose 12% of keys among the 10K hottest keys with a uniform distribution. We store the chosen keys as a text file to make experimental results consistent. The controller loads the file and puts the keys to the switch cache.
@@
5.2 Main Results
Throughput with different key access distributions. We measure the throughput with different skewness and plot the results in Figure 9. StarCache (switch) is the throughput of the cache in the switch. In other words, StarCache (servers+switch) is the total throughput, the sum of server throughput and switch throughput. We can see that StarCache provides high throughput regardless of skewness, unlike the others whose throughput decreases as the workload becomes more skewed. For example, StarCache provides high throughput even for highly skewed workloads with Zipf-0.99. NetCache does not provide high throughput as expected since many hot items are not cacheable. In the Zipf-0.99 workload, StarCache has higher throughput than NoCache and NetCache by 1.46$\times$ and 1.39$\times$, respectively.
Throughput on individual servers. We plot the load on individual servers in Figure 10. We can see that NoCache and NetCache do not balance loads well. This is due to the lack of caching logic for NoCache and many uncacheable hot items for NetCache. However, StarCache can balance the loads between the servers sufficiently since StarCache can cache variable-length items.
Latency. We measure the latency by varying target throughput. Figure 11 shows the median and the 99th percentile latencies as a function of throughput. Since there is no cache logic, NoCache does not provide low latency with limited throughput. NetCache is better than NoCache, but still does not provide the best performance. On the other hand, StarCache provides low latency in both the median and 99th percentile latencies while providing the highest throughput.
Impact of write ratios. To investigate the performance of StarCache in various workloads, we measure the throughput by varying the write ratios. Figure 12 reports the throughput as a function of write ratios. As expected, the throughput of StarCache decreases as the write ratio grows. This is because the switch invalidates the cached key when handling a write request for a cached item to ensure cache coherence. The switch cache provides benefits for only read requests, the throughput is decreased with higher write ratios since read requests for invalid cached items are forwarded to storage servers. With the 100% write ratio, the throughput of StarCache converges to NoCache since the cache does not provide any benefit. Meanwhile, similar to StarCache, NetCache offers reduced throughput as the write ratio grows since it also invalidates the cached key if there is a write for the key.
@@ 5.2 Main Results
Figure 11: Throughput vs. latency.
Figure 12: Impact of write ratios.
Impact of item size. Figure 13 shows the throughput with different key sizes and value sizes. In Figure 13 (a), we can see that the throughput of StarCache decreases as the key size grows. This is because larger keys consume more resources in storage servers and the switch. The latency to recirculate and process a cache packet especially increases, though it still remains at a few hundred nanoseconds. NoCache and NetCache also show similar results. The results with different value sizes in Figure 13 (b) are similar to those with different key sizes. StarCache shows the best performance except in a case with 64-B values. NetCache provides the highest throughput for the case since all items can be cacheable.
(a) Impact of key size
(b) Impact of value size
Figure 13: Impact of item size.
Figure 14: Impact of number of servers.
@@ 5.3 Deep Dive
Impact of number of servers. Figure 14 shows the individual throughput of 9, 18, and 36 storage servers. We also present the result with 36 servers for the uniform workload as the baseline. We can see that StarCache can balance loads sufficiently regardless of the number of servers. StarCache does not achieve perfect load balancing. This is because there are trade-offs to achieve variable-length caching, like limited cache size and extra latency caused by packet recirculation.
Impact of cache size. Figure 15 plots the throughput breakdown of StarCache with different cache sizes and the corresponding overflow request ratios. We can see that the total throughput increases as the number of cached entries grows. However, the throughput is saturated around 16-32 cached items. The throughput of the switch cache is also saturated as well. Figure 15 (b) clarifies the reason. From 64 cached items, the overflow request ratio rapidly increases, where the overflow requests indicate the requests for cached items forwarded to the storage servers due to the lack of free slots in the request table. Although the request table can maintain 256 request metadata for each key simultaneously, the slots get full quickly if the queueing between cache packets occurs due to too many cache packets.
(a) Throughput
(b) Overflow request ratio
Figure 15: Impact of cache size.
@@
Latency breakdown. Figure 16 (a) and (b) plot the median and the 99th percentile latency breakdowns, respectively. For the median latency, we can see that there is no distinct difference between NetCache and StarCache for servers. However, for StarCache, the latency of requests handled by the switch is slightly higher than NetCache by $1\mu s$. This is because StarCache requires requests to wait until circulating cache packets handle them. The difference in the caching logic between NetCache and StarCache results in more different tail latency as throughput grows. In Figure 16 (b), we can see that the tail latency of requests handled by the switch cache in StarCache gradually increases as throughput grows, unlike NetCache showing a deterministic latency even if throughput is saturated. This is because the queueing delay between requests for the same cached item occurs when the throughput gets closer to the saturated throughput.
Figure 17: Impact of multi-packet items.
Impact of multi-packet items. Figure 17 shows the performance of StarCache by considering that all cached items are multi-packet items. For one packet, we use a 1416-B value. For example, a 4-packet item has 16-B key and 5664-B values. We can see that StarCache generally balances loads even with multi-packet items. We observe that the balancing efficiency slightly decreases as the value sizes grow. This is because we decrease the cache size by considering that multiple cache packets take up the cache for a single item.
Impact of cacheable item ratios for NetCache. Figure 18 shows the throughput and the balancing efficiency of NetCache and StarCache with different ratios of 64-B values. This is to inspect whether NetCache is better than StarCache when it can cache more items. The balancing efficiency is defined as $LowSrv/HighSrv$ where $LowSrv$ is the storage server where the throughput is the lowest between the storage servers, and $HighSrv$ is the storage server with the highest throughput (i.e., the server with the hottest item). We use a near-optimal cache size for StarCache. The number of cached items is 32 for the ratios between 0-60%. We use 64 cache entries for 70-90%. For the ratio of 100%, we cache 128 hot items. Figure 18 (a) shows mixed results. StarCache outperforms until 50%, but is beaten by NetCache after 60%. We note that the throughput of NetCache in 80-100% is limited by the Tx throughput. However, NetCache does not balance loads sufficiently. In Figure 18 (b), we see that NetCache has poor balancing efficiency even if only 10% items are uncacheable. Unlike NetCache, StarCache can balance loads across all the ratios. We recall that the number of workloads where the portion of cacheable items for NetCache is higher than 50% is only 2 between 54 workloads [13].
@@
Figure 18: Impact of cacheable item ratios for NetCache.
Handling dynamic workloads. We investigate the performance of StarCache with dynamic workloads. We use a hot-in pattern, which is the most radical workload change. Specifically, every 10 seconds, the popularity of the 256 coldest items and the 256 hottest items is swapped. This means that all cache entries should be replaced. These sudden popularity changes are not common in practice, but this enables us to evaluate the robustness of StarCache to dynamic workloads.
Figure 19 (a) shows the throughput during 60 seconds. We can see that the throughput decreases when item popularity changes but recovers within a few seconds. This is because the controller in the switch control plane quickly updates the cache entries and fetches cache packets based on the top-$k$ report of servers and the cache popularity counter of the switch data plane. Figure 19 (b) plots the change in the cache size over time with a given threshold of 0.05% and a minimum cache size of 32. The controller adjusts the cache size dynamically by monitoring the overflow request ratio. We can see that the cache size fluctuates between 32 and 36, hence the switch can maintain the overflow request ratio within 0.05%.
@@ 6 Related work
In-network caching. We have witnessed a surge of domain-specific acceleration through in-network computing (e.g., consensus [15], coordination [21], machine learning [35], replication [26]). Storage systems are major target domains of recent works (e.g.,  [22, 26, 39, 24, 25, 40, 28, 20, 29, 19, 12, 23]). NetCache [22] is an early work that has a great impact not only on in-network caching but also on the entire field of in-network computing. It shows the potential and limitations of the programmable switch as an in-network accelerator for distributed systems. DistCache [33] resolves inter-rack load imbalance and FarReach [36] realizes write-back caching, but they can still cache a limited size of items. StarCache enables variable-length in-network caching by leveraging built-in features of the programmable switch while sacrificing the cache size reasonably.
Other alternatives. SwitchKV [30] employs the OpenFlow switch for cache lookups while items are cached in cache nodes, which provide limited performance. Pegasus [28] is a selective replication solution where the switch replicates hot items across storage servers for load balancing, functioning similarly to caching. Unfortunately, its performance is limited by the aggregate throughput of storage servers. StarCache provides high-performance caching beyond the aggregate throughput with no extra hardware costs.
@@ 7 Conclusion
We proposed StarCache, an in-network caching architecture that is capable of variable-length caching in programmable switches. The key idea of StarCache is to make cached key-value pairs revisit the switch data plane continuously by leveraging packet replication efficiently. We addressed technical challenges to realize the idea of StarCache. We have implemented a StarCache prototype on an Intel Tofino switch, and experimental results demonstrated that StarCache can balance skewed workloads with diverse sizes of items. We believe that StarCache is a fine example of overcoming fundamental switch hardware constraints that put limits on application semantics, which are major hurdles of in-network computing research.
Acknowledgement
The author thanks Yeon-sup Lim for providing constructive feedback on the earlier draft of the paper.

### v3

@@ 4 Implementation
Client-server application. We develop an open-loop application in C using NVIDIA Messaging Accelerator library (VMA) [2]. VMA bypasses kernel network stacks by intercepting the socket function calls and translating them to native RDMA verbs. The client application measures throughput and latency by generating requests. The time gap between consecutive requests follows an exponential distribution. The server application has multiple threads where each thread is pinned to a disjoint CPU core. To emulate multiple storage servers, we assign a partition per thread so that each thread acts as an independent storage server. We also limit the Rx throughput of each emulated server to 100K RPS to ensure the bottleneck is at servers in our testbed. This technique is used in existing works [21, 29, 34] as well. In a similar vein, like NetCache [21], we implement a key-value store with TommyDS [1], a high-performance hash table library.
Switch. We implement the switch data plane in P416 [9] for Intel Tofino [5]. We use Intel P4 Studio SDE 9.7.0 to compile the switch data plane. Our prototype uses 9 stages and 6.67% SRAM, 7.38% Match Input Crossbar, 9.29% Hash Bit, and 30.56% ALUs. The request table has a maximum queue size of 8 for each key. The controller is written in Python 3. In our prototype, we use an additional register array for the request table to store the timestamp of the request for latency measurement. The OrbitCache header has 3 extra fields of 1-B Cached, 4-B Latency, and 1-B SrvID. The first two fields are required to separately measure the latency of requests for cached and uncached items. The last field is to store the emulated server ID as we emulate multiple storage servers as dedicated threads in a physical node.
@@ 5.1 Methodology
Testbed setup. We build a cluster consisting of 8 nodes, which are connected by an APS BF6064X-T switch with the Intel Tofino 1 [5]. The servers are equipped with a 10-core CPU (Intel i5-12600K @ 3.7 Ghz, 12 hyperthreads and 4 non-hyperthreads), 32 GB of DDR5 memory, and a 100GbE NVIDIA CX-5 NIC. The servers run Ubuntu 22.04 LTS with Linux kernel 6.5.0. The 4 nodes act as clients, and the remaining 4 nodes are used to emulate multiple storage servers.
@@
Workloads. We emulate a single storage rack with 32 storage servers using 8 partitioned threads per node. We basically consider a workload with 10M key-value pairs whose key popularity skewness follows a Zipfian distribution with $\alpha=0.99$, since it is regarded as typical skewness [13, 7]. Although both key and value sizes impact whether an item is cacheable by NetCache, we use 16-byte keys by default for simplicity, the maximum supported key size by NetCache. Instead, we represent the portion of cacheable items using different ratios between 64 bytes and 1024 bytes values. The 64-byte value represents a cacheable item value of NetCache. The 1024-byte value is a typical-sized item value in many workloads and is considered in many research papers [29, 30, 31, 17, 6, 26]. We use a bimodal distribution with 82% 64-byte and 18% 1024-byte values by considering the cacheable item ratio of NetCache for the Cluster018 workload of Twitter [12]. Most experiments are for read-only workloads as we target read-intensive workloads.
Except for dynamic workloads, we preload the 10K and 128 hottest items for NetCache and OrbitCache, respectively. 128 is a nearly optimal cache size for OrbitCache that provides the best performance gains. Since NetCache can cache only 82% of items in the workload, the actual number of cached items is 8200, which is still larger than the cache size of OrbitCache by 64$\times$. To be fair, we choose 82% of keys among the 10K hottest keys with a uniform distribution. We store the chosen keys as a text file to make experimental results consistent. The controller loads the file and puts the keys to the switch cache. We note that, even if we use different key samples, trends have been consistent.
5.2 Main Results
Throughput with different key access distributions. We measure the throughput with different skewness and plot the results in Figure 9. We can see that OrbitCache provides high throughput regardless of skewness, unlike the others whose throughput decreases as the workload becomes more skewed. NetCache does not provide high throughput as expected since many hot items are not cacheable. In the Zipf-0.99 workload, OrbitCache has higher throughput than NoCache and NetCache by 3.59$\times$ and 1.95$\times$, respectively. The server throughput in OrbitCache is consistent across the given skewness, and this means that the loads are balanced.
@@ 5.2 Main Results
Individual server loads. We plot the load on individual servers in Figure 10. We can see that NoCache and NetCache do not balance loads well. This is due to the lack of caching logic for NoCache and many uncacheable hot items for NetCache. However, OrbitCache can balance the loads since OrbitCache can cache variable-length items.
Latency vs. throughput. We measure the latency by varying Tx throughput. Figure 11 shows the median and the 99th percentile latencies as a function of Rx throughput. OrbitCache provides the best throughput but slightly higher latency than NetCache of 1 microsecond. This is because NetCache has a larger cache size, and requests handled by the switch dominate in latency data. In addition, in OrbitCache, requests for cached items should wait until cache packets read them. Although NetCache has slightly better latency, throughput is very limited since it fails to balance loads.
Impact of write ratio. Figure 12 reports the throughput as a function of write ratios. The throughput of OrbitCache decreases as the write ratio grows. This is because the switch invalidates the cached key when handling a write request for a cached item to ensure cache coherence. The switch cache provides performance gains for read requests, the throughput is decreased with higher write ratios since read requests for invalid cached items are forwarded to storage servers. With the 100% write ratio, the throughput of OrbitCache converges to NoCache since the cache does not provide any benefit. Meanwhile, similar to OrbitCache, NetCache offers reduced throughput as the write ratio grows since it also invalidates the cached key if there is a write for the key.
@@
Figure 11: Latency vs. throughput.
Figure 12: Impact of write ratio.
(a) Throughput
(b) Balancing efficiency
Figure 13: Scalability.
Scalability. We now inspect whether OrbitCache can balance loads of different numbers of servers. In this experiment, we limit the Rx throughput to 50K RPS to ensure that the bottleneck occurs at the storage servers rather than the clients, even when using 64 servers. In Figure 13 (a), we can see that the throughput of OrbitCache is improved almost linearly while the others do not. Figure 13 (b) clarifies that this is because NoCache and NetCache fail to balance imbalanced loads. Balancing efficiency is defined as the minimum throughput between the servers divided by the maximum throughput between the servers.
We note that OrbitCache may not balance workloads with larger servers like hundreds of servers because the cache size is limited. OrbitCache targets a single storage rack consisting of tens of servers, and our experiment result demonstrates that OrbitCache is enough to balance a rack-scale storage system. To scale out for hundreds of servers, instead of relying on a single switch to balance all servers, each ToR switch in clusters should cache items of storage servers belonging to its rack, as described in Section 3.9.
Performance with production workloads. We conduct experiments with several workloads of Twitter [37] to see how OrbitCache works with various production workloads. We pick 5 workloads based on the cacheable item ratio. We assign the workload ID A to E for Cluster045/016/044/017/020. We still use the 16-B keys for simplicity but vary the write ratio, the portion of 64-B values. Unlike the other experiments, the cacheable item ratio is controlled by choosing keys with a uniform distribution independent of the portion of small 64-B values.
Figure 14 shows the results. Although the difference in throughput varies depending on the used workloads, we can see that OrbitCache shows the best performance for all the workloads. There is a little difference for Workload A. This is because NetCache can cache 95% of items, and the write ratio is relatively high. Workload E is a workload that shows a significant performance gap. This is because only 1% of items are cacheable for NetCache.
Figure 14: Performance with production workloads.
@@ 5.3 Deep Dive
Latency breakdown. Figure 15 (a) and (b) plot the median and the 99th percentile latency breakdowns, respectively. In the median latency, OrbitCache shows slightly higher switch latency than NetCache. This is because requests in OrbitCache should wait until circulating cache packets read them, resulting in latency overhead. In Figure 15 (b), we can see that the tail latency of OrbitCache (switch) increases as throughput grows. This is not surprising because we use a circular queue-based request table to maintain request metadata and packet cloning for a cache hit, resulting in additional latency overhead. We believe that this is acceptable since OrbitCache can provide much higher throughput than the others. In addition, the tail latency of the switch is still tens of microseconds even when the tail latency of servers soars by reaching the saturated throughput. We can decrease the tail latency of the switch by reducing the cache size as well. If the cache is based on a commodity server, the tail latency would be 10-100$\times$ longer than the switch cache.
Impact of cache size. Figure 16 plots the throughput breakdown, request latency handled by the switch, and the overflow request ratio of OrbitCache with different cache sizes. We can see that the total throughput increases as the number of cached entries grows. However, the throughput is saturated with around 128 cached items. Similarly, in Figure 16 (b), the tail latency increases quickly after 64-128 cached items. Figure 16 (c) clarifies the reason. From 256 cached items, the overflow request ratio rapidly increases, where the overflow requests indicate the requests for cached items forwarded to the storage servers due to the lack of free slots in the request table. The lack of free slots is caused by the excessive queueing delay between too many cache packets beyond the capability of the switch hardware. This clarifies that the cache size is the key to determining the performance. This also means we should choose an effective cache size between 32 and 128 to balance the trade-off.
(a) Saturated Throughput
(b) Switch Cache Latency
(c) Overflow request ratio
Figure 16: Impact of cache size.
@@
(c) Cache size
Figure 17: Impact of item size.
Impact of item size. We measure throughput, balancing efficiency, and the maximum cache size by varying the item size. In this experiment, we use the 100% same value size for all items to inspect how OrbitCache works in the worst case. Note that 16-B key and 1416-B value are the maximum size for a single packet payload with 28-B custom header fields.
Figure 17 (a) shows that OrbitCache can balance various workloads, including a workload where 100% items are MTU-sized. The slight drops in throughput are due to the increased value size. Figure 17 (b) shows the balancing efficiency. We see that OrbitCache generally maintains high balancing efficiency. Figure 17 (c) plots the effective cache size that provides the highest performance gain. We can see that the effective cache size decreases as the value size increases. This is because larger cache packets consume more switch resources, making the effective cache size small. Nevertheless, we can see that OrbitCache can balance the workloads well since we need only a small number of cached items to balance skewed workloads. Due to the page limit, we omit the result in the impact of key size, which is similar to the result in the impact of value size.
Handling dynamic workloads. We investigate how OrbitCache reacts to dynamic workloads. Like existing works [21, 34, 29], we use a hot-in pattern, which is the most radical workload change. We use 4 storage servers without server emulation and Rx rate limits similar to a previous work [34] to avoid inaccuracy due to the system state change. Here, every 10 seconds, the popularity of the 128 coldest items and the 128 hottest items is swapped. This means that all cache entries should be replaced suddenly. These sudden popularity changes may not be common in practice, but this enables us to evaluate the robustness of OrbitCache.
Figure 18 (a) shows the throughput for 60 seconds. We can see that the throughput decreases when key popularity changes but recovers within a few seconds. This is because the controller in the switch control plane quickly updates the cache entries and fetches cache packets based on the top-$k$ report of servers and the cache popularity counter of the switch data plane. Figure 18 (b) plots the change in the overflow request ratio over time. We can see that the ratio soars when the key popularity changes. This is because it takes time to fetch items from storage servers. Similar to the throughput result, the overflow ratio decreases after the items are fetched. It would be interesting if we could design a mechanism that minimizes the cache update time.
@@ 6 Related work
In-network caching. Storage systems are major target domains of recent in-network computing works (e.g.,  [21, 25, 39, 23, 24, 40, 27, 19, 28, 18, 11, 22, 36, 38, 20, 14]). NetCache [21] shows the potential and limitations of switch-based caching. The limitation in the item size is passed down to recent works as well. FarReach [34] enables write-back caching for NetCache. DistCache [32] proposes a distributed caching architecture using power-of-two choices and hash-based cache allocation to balance multiple clusters. These solutions are based on the NetCache architecture, hence they can still cache a limited size of items. OrbitCache enables variable-length in-network caching by leveraging built-in features of the switch.
Other alternatives. SwitchKV [29] employs the Openflow switch for cache lookups while items are cached in a cache node. Although this reduces the lookup overhead, it still provides limited performance due to server-based caching. Pegasus [27] is a selective replication solution where the switch replicates hot items across storage servers for load balancing. Unfortunately, its performance is limited by the aggregate throughput of storage servers. OrbitCache provides high-performance single-box caching beyond the aggregate throughput of commodity storage servers.
@@ 7 Conclusion
We proposed OrbitCache, an in-network caching architecture that is capable of variable-length caching in programmable switches. The key idea of OrbitCache is to make cached key-value pairs revisit the switch data plane continuously by leveraging packet replication efficiently. We have implemented a OrbitCache prototype, and experimental results demonstrated that OrbitCache can balance skewed workloads with diverse conditions. We hope that this work can contribute to the research community by providing insights that utilizing built-in hardware features has great potential to make in-network computing solutions effective.